A feather is dropped on the moon from a height of 1.40 meters. The acceleration of gravity on the moon is 1.67 m/s^2. Determine the time for the feather to fall to the surface of the moon. *


1
Expert's answer
2022-02-13T12:14:03-0500

Given:

"h=1.40\rm\: m"

"g=1.67\:\rm m/s^2"


The time required to get a surface of Moon by the feather

"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2*1.40}{1.67}}=1.29\:\rm s"